a. The sales representative here tells you they also have two floor plans available, but they only have 38 homes still for sale. Write an equation that illustrates the situation. Use x and y to denote floor plan one and floor plan two respectively.
b. The representative tells you that floor plan one sells for $175,000 and floor plan two sells for $200,000. She also mentions that all the available houses combined are worth $7,200,000. Write an equation that illustrates this situation. Use the same variables you used in Part a.
c. Use elimination to determine how many houses are available in each floor plan. Explain how you arrived at your answer.
I know part A is X+Y=38 however, I am no good at the elimination method

Answer by AKT(2)

x^2-6x-16=0
then
x^2-8x+2x-16=0
then
x(x-8)+2(x-8)=0,
(x-8) (x+2) =0,
therefor
x=8 or x=-2.
-2 is not a possible answer therefor x=8 is the final answer.